OpenStudy (anonymous):

r/12=1 9/10 PLZ HELP WILL FAN AND MEDAL what is r?

OpenStudy (anonymous):

@SolomonZelman and @Godlovesme

OpenStudy (anonymous):

oh and @justuu

OpenStudy (anonymous):

and @TheSmartOne

OpenStudy (anonymous):

Just tell me if I'm right or wrong.

OpenStudy (anonymous):

Is it 24 4/5

OpenStudy (anonymous):

@TheSmartOne please help!!!

OpenStudy (anonymous):

@justuu, @SolomonZelman and @Godlovesme

OpenStudy (anonymous):

@lizz123 plz help I just chose a random person!!

OpenStudy (godlovesme):

@kdenice123 hold on i'm checking

OpenStudy (godlovesme):

first change it to improper fraction

OpenStudy (anonymous):

I think I got the answer, just tell me if i am right or wrong cuz I have another thing in mind.

OpenStudy (anonymous):

24 4/5?

OpenStudy (godlovesme):

i didn't get that i got 22.8

OpenStudy (anonymous):

Well the answers in my book do not have that.

OpenStudy (godlovesme):

\[r/12= 1 9/10 \] 1 9/10 ***

OpenStudy (lizz123):

wht choices do u have

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@SolomonZelman HELP PLEASE!!!!

OpenStudy (godlovesme):

1 9/10 ==>> 19/10 right @kdenice123

OpenStudy (anonymous):

I knew it!!!

OpenStudy (godlovesme):

\[r/12=19/10\] then u can cross multiply

OpenStudy (godlovesme):

\[10r=288\] \[r=28.8 \]